OpenStudy (anonymous):

The base of a triangle is 3cm ,greater then the height the area is 44cm^2 find the height and the length of the base

OpenStudy (anonymous):

h=height 3+h=base 1/2h(3+h)=44

OpenStudy (anonymous):

\[h^2+3h=88\]\[h^2+3h−88=0\]

OpenStudy (anonymous):

\[(h+11)(h-8)=0\]

OpenStudy (anonymous):

h=8cm

OpenStudy (anonymous):

3+h=b ~ 3+8=b b=11

OpenStudy (anonymous):

and that's it :)